The European Currency Symbol for L^{A}TEX

by Henrik Theiling

                         theiling@coli.uni-sb.de

Why?

The European currency symbol is already available for L^{A}TEX in different packages (Text-Companion fonts, Marvosym package, etc.). However, I wanted to create a symbol that is constructed according to the official European Commision’s definitions. Furthermore, I wanted to do it with METAFONT because I don’t like to use PostScript fonts because they are likely to create compatibility problems.

Usage

At the beginning of the document in the pre-amble, declare \usepackage{eurosym}. Then the new com-

mands \officialeuro{}, \geneuro{}, \geneuronarrow{} and \geneurowide{} create , C, C, and C

resp. The latter three create an overlayed symbol from the current font’s C and the two horizontal bars with three different lengths. Of course you should only use the latter commands if the font you’re using lacks the symbol or you don’t like “ ” for some reason. Officially, “ ” has to be used with all the fonts because it’s the only official shape. However, this doesn’t always look nice (especially in bold or oblique font shapes). Furthermore, even the OCR draft suggests a different shape for OCR-B.

There is also the command \euro{} which defaults to be a shortcut for \officialeuro{}. You can set a different default symbol by either declaring e. g. \let\euro=\eurogen or by using one of the package

options official, gen, gennarrow or genwide.

There is a convenient command \EUR{… } which lets you typeset an amount of money nicely (with a micro space \, between the symbol and the number). Because in some countries the symbol has to appear on the left of the number while in others it has to be on the right, the packages recognizes the two options left and right. So if you put \usepackage[left]{eurosym} at the beginning of your document, \EUR{1000} will create 1000, while a \usepackage[right]{eurosym} makes it appear as 1000 . The package default is [left] unless the german package was included before the eurosym package. You can change the shape of the symbol that \EUR uses by redefining \euro.

Table of Commands

Here is a table of the major commands:

\usepackage[options]{eurosym} include the eurosym package. Available options:

                               left, right, official, gen, gennarrow, genwide.

\euro{}  create a symbol. The shape depends on the

package options and defaults to \officialeuro{}

\EUR{amount}  typeset an amount of . The position of the

currency symbol depends on the package option and defaults to left except the german package is loaded in which case it defaults to right.

There should generally be no need to use the following minor commands.

\officialeuro{}  create a symbol

\geneuro{}  create a C symbol

\geneuronarrow{} create a C symbol
\geneurowide{}  create a C symbol

\eurobars{}  create only the bars of the symbol:
\eurobarsnarrow{} create the bars of the symbol in 80% width:
\eurobarswide{}  create the bars of the symbol in 120% width:

Exact Sizes

A symbol is as tall as a C. The bx-shaped version should be a little wider than the normal one and should of course be bold.

24.88 pt: C 10 pt: C 6 pt: C  n
17.28 pt: C  9 pt: C  5 pt: C  b
14.4 pt: C  8 pt: C
12 pt: C  7 pt: C

Appearance in Text

m n Ich bezahlte 500 für das Radio. Bzw. 1000 C für den Fernseher.
m sl Ich bezahlte 500 für das Radio. Bzw. 1000 C für den Fernseher.
m it Ich bezahlte 500 für das Radio. Bzw. 1000 C für den Fernseher.
m sc Ich bezahlte 500 für das Radio. Bzw. 1000 C für den Fernseher.
bx n Ich bezahlte 500 für das Radio. Bzw. 1000 C für den Fernseher.
bx sl Ich bezahlte 500 für das Radio. Bzw. 1000 C für den Fernseher.
bx it Ich bezahlte 500 für das Radio. Bzw. 1000 C für den Fernseher.

Table of Shapes

The following shapes are derived from the official symbol “Euro glyph”.

    n=sc  sl=it  ol
m
bx

The style file defines \slshape as \itshape for this symbol and normal shape for \scshape.

Table of Generic Shapes

The font also contains only the bars for a fast hacking way to create the Euro currency symbol with fonts that don’t contain it. Usually you can simply use \geneuro to get a hacked Euro symbol for the current font.

n   sl m C C bx C C

If the font you are using is wider or more narrow so that the sizes of the bars don’t look nice for that font,

you can either try \geneuronarrow or \geneurowide.

     \geneuronarrow \geneurowide
      n       sl       n      sl

m C  C  C  C
bx C  C  C  C

Construction of the Symbol

The construction is taken from the German c’t Magazine, 11/98, page 211. Let the line thickness be x. Then the radius of the inner circle is 5 x and the distance between the inner bars is x. The angle of the opening on the left is 80^{∘}. All the other points are obtained by intersection of lines. However, I don’t know how to construct the amount the bars leave the symbol on the left (marked with a). If someone knows how it’s really done, please tell me. I constructed them my using a ruler on a sample I had, so it should be quite exact.

80^{∘}^{⟵ a = ? ⟶}

Example in a Longer Text

In the following, I’ve copied an article from a local newspaper (Neue Westfälische, Nr. 174, Donnerstag,

30. Juli 1998) containing money amounts and changed “DM” to “ ” or “C” resp. in order to give an impression of how it looks in a longer text.

Erzeugergemeinschaft plant bis zum Jahr 2003 Verdopplung des Umsatzes

EGO will Riesenvorsprung nutzen

B i s s e n d o r f / L a g e (blo). Verbraucher kaufen Fleischund Wurstwaren inzwischen sehr kritisch ein, gehen wieder viel häufiger ins Fleischerfachgeschäft. Das kommt der EGO (Erzeugergemeinschaft für Schlachtvieh im Raum Osnabrück e. G.) mit Ihren ”Eichenhof“- Produkten entgegen. Die EGO setzt auf nachprüfbare Herkunft und Qualität, kooperiert mit 175 Fleischereien und plant bis zum Jahr 2003 eine Umsatzverdopplung auf 400 Mio. .

Dabei kann die EGO ein großes Pfund in die Waagschale werfen. ”Wir haben mindestens 15 Jahre Vorsprung.“ erklärte geschäftsführender Vorstandsvorsitzender Karl-Heinz Hügelsmeyer in Bissendort. Der Vorsprung sind die strengen Kriterien, nach denen der genossenschaftliche Zusammenschluß vor knapp 700 vertraglich gebundenen bäuerlichen Familienbetrieben zwischen Teutoburger Wald und Wiehengebirge arbeitet: Tiergerechte Schweineund Rinderhaltung mit festen Regeln für Fütterung, Zucht und Aufzucht, Andienungspflicht und Abnahmegarantie, Sauberkeit der Produktion, Regelverstöße werden hart mit Auschluß geahndet.

Die Landwirte profitieren durch gute Auszahlungspreise und Prämien an die Mitglieder in 1997 aus, berichtete Geschäftsführer Rudolf Fester. Er löst am 1. August EGO-Gründer Karl-Heinz Hüggelsmeyer als Vorstandschef ab. Der 65jährige Hüggelsmeyer wird noch für einige Jahre als Geschäftsführer der Tochterfirmen Pieper (Lage) und Kinnius (Osnabrück) tätig sein. diese beiden Verarbeitsbetriebe erzielen den Angaben zufolge derzeit positivere Ergebnisse als 1997, weil die Rohstoffpreise sinken. Insgesamt stehe die EGO besser da als vor einem Jahr. Für 1998 rechnet die 210 Mitarbeiter beschäftigende Gruppe mit 200 Mio. , davon 5 Mio. mit Convenience-Produkten. Sie sollen einmal 15 Mio. bringen. 1997 war der EGO-Umsatz um 7% auf 197 Mio. gestiegen.

Die Landschlachterei Pieper, die einige ihrer Abnehmer ausgesiebt hat, kam dabei im Vorjahr auf 29,3 (Vorjahr 30,3) Mio. . Die Zahl der Mitgliedsbetriebe stieg um gut 100 auf 687. Aus der Fusion mit der Erzeugergemeinschaft Minden-Ravensberg-Lippe (Herford), die 320 Mitglieder hatte, kamen nur 81 Betriebe hinzu. Das Gros wurde nicht übernommen. Zitat: ”Die wollten unsere Kriterien nicht erfüllen.“

Erzeugergemeinschaft plant bis zum Jahr 2003 Verdopplung des Umsatzes

EGO will Riesenvorsprung nutzen

B i s s e n d o r f / L a g e (blo). Verbraucher kaufen Fleischund Wurstwaren inzwischen sehr kritisch ein, gehen wieder viel häufiger ins Fleischerfachgeschäft. Das kommt der EGO (Erzeugergemeinschaft für Schlachtvieh im Raum Osnabrück e. G.) mit Ihren ”Eichenhof“- Produkten entgegen. Die EGO setzt auf nachprüfbare Herkunft und Qualität, kooperiert mit 175 Fleischereien und plant bis zum Jahr 2003 eine Umsatzverdopplung auf 400 Mio. C.

Dabei kann die EGO ein großes Pfund in die Waagschale werfen. ”Wir haben mindestens 15 Jahre Vorsprung.“ erklärte geschäftsführender Vorstandsvorsitzender Karl-Heinz Hügelsmeyer in Bissendort. Der Vorsprung sind die strengen Kriterien, nach denen der genossenschaftliche Zusammenschluß vor knapp 700 vertraglich gebundenen bäuerlichen Familienbetrieben zwischen Teutoburger Wald und Wiehengebirge arbeitet: Tiergerechte Schweineund Rinderhaltung mit festen Regeln für Fütterung, Zucht und Aufzucht, Andienungspflicht und Abnahmegarantie, Sauberkeit der Produktion, Regelverstöße werden hart mit Auschluß geahndet.

Die Landwirte profitieren durch gute Auszahlungspreise und Prämien an die Mitglieder in 1997 aus, berichtete Geschäftsführer Rudolf Fester. Er löst am 1. August EGO-Gründer Karl-Heinz Hüggelsmeyer als Vorstandschef ab. Der 65jährige Hüggelsmeyer wird noch für einige Jahre als Geschäftsführer der Tochterfirmen Pieper (Lage) und Kinnius (Osnabrück) tätig sein. diese beiden Verarbeitsbetriebe erzielen den Angaben zufolge derzeit positivere Ergebnisse als 1997, weil die Rohstoffpreise sinken. Insgesamt stehe die EGO besser da als vor einem Jahr. Für 1998 rechnet die 210 Mitarbeiter beschäftigende Gruppe mit 200 Mio. C, davon 5 Mio. C mit Convenience-Produkten. Sie sollen einmal 15 Mio. C bringen. 1997 war der EGO-Umsatz um 7% auf 197 Mio. C gestiegen.

Die Landschlachterei Pieper, die einige ihrer Abnehmer ausgesiebt hat, kam dabei im Vorjahr auf 29,3 (Vorjahr 30,3) Mio. C. Die Zahl der Mitgliedsbetriebe stieg um gut 100 auf 687. Aus der Fusion mit der Erzeugergemeinschaft Minden-Ravensberg-Lippe (Herford), die 320 Mitglieder hatte, kamen nur 81 Betriebe hinzu. Das Gros wurde nicht übernommen. Zitat: ”Die wollten unsere Kriterien nicht erfüllen.“